Find area bounded by parabola y2 = 4ax and line y = mx.


1
Expert's answer
2021-05-02T08:08:36-0400

The area enclosed between the parabola "y^2=4ax" , and the line "y=mx" , is represented by the shaded area"OABO" as



The points of intersection of both the curves are"(0,0)" and"(\dfrac{4a}{m^2},\dfrac{4a}{m} )" .

We draw AC perpendicular to "x-axis" .

So, "Area OABO= Area OCABO - Area (OCA)"

"= \int_0^{\dfrac{4a}{m^2}}2\sqrt{ax}-" "\int_0^{\dfrac{4a}{m^2}}mxdx"


="2\sqrt{a}[\dfrac{x^{1.5}}{1.5}]_0^\dfrac{4a}{m^2}" "-" "m[\dfrac{x^2}{2}]^{\dfrac{4a}{m^2}}_0"


"=\dfrac{8a^2}{3m^3} units"